A particle moves with constant acceleration for 6 seconds after starting from rest. The distance travelled during the consecutive 2 seconds interval is in the ratio?


1
Expert's answer
2021-09-02T10:42:16-0400
"s_1=0.5a(1)^2=0.5a\\s_2=0.5a(2)^2-0.5a(1)^2=1.5a\\s_3=0.5a(3)^2-0.5a(2)^2=2.5a\\s_1:s_2:s_3=1:3:5"